OpenStudy (anonymous):

-4x+3y evalute each expression for x=3 and y=2

OpenStudy (anonymous):

plug in x and y values is how you solve this I believe.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-4%283%29%2B3%282%29 result = -6

OpenStudy (anonymous):

can u atleast show me the steps

jigglypuff314 (jigglypuff314):

for -4x+3y when x = 3 and y = 2 you plug in what x equals for x and what y equals for y so -4(3) + 3(2) then you multiply and add

jigglypuff314 (jigglypuff314):

@HelpMePlease? as much as I love wolframalpha as well, it is best to try to explain to someone how to do the problem in the comments section of the problem, so that it would be visually easier to understand :)

OpenStudy (anonymous):

i nee help on this question four times the sum of a number and -3 is 4 more than twice the number write and solve an equation to find the number. please help and solve it 4 me plz

jigglypuff314 (jigglypuff314):

4 * (x + (-3)) = 4 + 2x I'm not good with word problems but is this what the question is asking?